RC Meeting, Tuesday, October 26 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920

Accelerator Status

  1. The accelerator is still recovering from the tour, the repair of magnet MYBBS04, and a storm that passed thru around midnite.
  2. No beam has been seen yet downstream of the repaired magnet, so no beam to Hall D yet.
  3. The pass changes for Halls A and B have taken place.
  4. As of 08:00, MCC was still tuning to reduce Hall C beam contributing to high halo rates in Hall B.
  5. I think we're going to see a little more beam in Hall C, then they'll take it away to enter the accelerator to check magnet polarities and BPMs.

Hall C Status

  1. As of 08:00, since the beam was turned off yesterday for the tour, we had less than 15 minutes of 42 muA beam in Hall C.
  2. Number of good pions is 5189. (The needle barely moved since yesterday's 8am number.)

Current Run Plan

  1. Day shift only: Try to take 1 hour of Dummy target each day. (a 100% live hour)
  2. Tentatively scheduled: a BCM calibration on Wednesday mid-afternoon.
  3. Continue data taking until early Nov

Current Issues

  1. Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
  2. Q3 insulating vacuum spikes generating alarms. For example, https://logbooks.jlab.org/entry/3930175 .
  3. One HB LHe incident of note: https://logbooks.jlab.org/entry/3929152 . (possible clue to the origin?)
  4. Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
  5. Still running with no UPS on the SHMS cryo controls - Steve L. has ordered batteries and UPS. ( PENDING )
  6. HCCAM02 not functioning (lift required to replace).

Special Requests

Useful Links